If the area of circle inscribed in an equilateral triangle is 154 square cm. Find the perimeter of triangle.
 

Grade:10

1 Answers

jitender
6 years ago
Area of circle=^r2=154^=22/7And r= 7cmIn an equilateral triangle the median is divided in the ratio 2:1 And median = cmWhere a = side of equilateral triangle and perimeter of triangle = 3a= 42 =14a/2Here a comes a=42/a/2Put x= 721=a/23x== under rootLet x= 7Here x comes x= 7And 2x +x = a/2